OpenStudy (anonymous):

A particle travels 16 times around a 18-cm radius circle in 37 seconds. What is the average velocity (in m/s) of the particle?

OpenStudy (anonymous):

oh jeeze. 2 different answers lol

OpenStudy (anonymous):

and they're gone...

OpenStudy (anonymous):

so T is period right? and a period is how long it takes for one revolution so 37/16=2.3125

OpenStudy (anonymous):

and V=2pir/T I think

OpenStudy (anonymous):

Average velocity is displacement over time. Since the particle ends in the same spot as it started its average velocity is 0. If you're looking for average speed it would be total distance traveled over time: \[D=(2π∗r)*16=2∗3.14∗18*16=1808cm=18.08m\] V=D/t=18.08/37=.489m/s

OpenStudy (anonymous):

so V=((2)(pi)(18))/2.3125

OpenStudy (anonymous):

am I wrong?

OpenStudy (anonymous):

V=48.9

OpenStudy (anonymous):

I just need confirmation that I'm doing it right lol

OpenStudy (anonymous):

I think I did it right, I just forgot to convert cm to m

OpenStudy (anonymous):

If you are finding average speed, yes you're doing it right (except check your units). Average velocity would be 0 m/s

OpenStudy (anonymous):

so the answer is .489 m/s?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[V=r \omega\] \[\omega =(16*2\pi)\div37 =2.7 radians/\sec\] \[r=0.18 m\] \[Velocity, V=0.18*2.7=0.486m/s\]